Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 655     Entry time: 08 Oct 2025, 09:45
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 08.10.2025 
General Announcement:
- M. Doering announces on the LIMP-25: 
1. Karabo deployment ongoing; 
2. PLC updates see works at FXE, HED for next week cw42 (SQS later time); 
3. Central gas management will be implemented by TS - Tobias Clever informs that since yesterday company is onsite for programming the system changes. For the next couple of days 
a short interruptions of gases bottles may happen and to be finished by the end of next week.
- On Monday M. Doering sent an e-mail regarding the next Winter maintenance "light" for recurring activities that always happens during a regular winter maintenance. 
If groups have new tasks (unknown by the planning phase of the LIMP) that shall be done prior the start of operation, please let M. Doering know, especially if may affect other 
groups. 
For example, emergency OFF will be triggered in SA1 and SA2 instruments as well as cooling water will be interrupted for one day. 

CTRL:
- Proceeding with Karabo updates.

DA:
- Nothing to report.

EEE:
- Timing system update happening at SA3 instruments today.

VAC:
- Not present.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- Everything fine; yesterday the work in pre-absorber was finished well.
- This week the pre-alignment of ODL in SA3 is planned to be finished and then Vacuum can be closed.

LAS:
- Nothing to report.

UNSYS:
- Nothing to report.

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- Not present.

SCS:
- Next week works on the hutch will be made by an external company (area to be fenced). Please contact Jan Torben if access is needed.

SQS:
- Not present.

SXP:
- Nothing to report.
ELOG V3.1.4-7c3fd00